OpenStudy (anonymous):

The acceleration due to gravity on the Moon is about one-sixth what it is on Earth. If an object is thrown vertically upward on the Moon, how many times higherwill it go than it would on Earth, assuming the same initial velocity?

OpenStudy (anonymous):

is it just 6?

OpenStudy (anonymous):

yes its just six consider v^2= u^2+2as here 0=u^2-2gh h=u^2/2g if g became g/6 h becomes 6 times